OpenStudy (anonymous):

I need to find a point that is a fraction of the distance between two other points, and I can't remember how to do it.

OpenStudy (jdoe0001):

2 ducks and 3 pelicans

OpenStudy (anonymous):

For sake of clarification I have to find a point that is 2/5 the distance between (6,2) and (8,4)

OpenStudy (jdoe0001):

ok so let's take a peek a that say we find the whole distance between those 2 points and we divide it in say 5 pieces so 2 pieces on one side.. and 3 pieces on the other side of that distance segment what ratio would that make from the (6,2) to (8,4)?

OpenStudy (anonymous):

What do you mean, what ratio can I get from that?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

How is yes an answer that question...?

OpenStudy (jdoe0001):

notice we need to find the 2/5 distance so we'd first split the distance in 5 pieces... and let us find the ratio of those 2 sides

OpenStudy (jdoe0001):

well.. I meant yes, that's what we're looking for, the ratio.... what ratio would that make? :)

OpenStudy (anonymous):

2 to 3?

OpenStudy (anonymous):

I could've sworn there was just some formula for this... That's all I need.

OpenStudy (jdoe0001):

yeap 2 to 3 or 2:3 FROM (6,2) to (8,4) or  so just gimme a few secs

OpenStudy (jdoe0001):

\(\bf A(6,2)\qquad B(8,4)\qquad ratio1=2\qquad ratio2=3\qquad 2:3\\ \quad \\ \quad \\ \cfrac{AB}{BC}=\cfrac{ratio1}{ratio2}\implies ratio2\cdot AB=ratio1\cdot BC\quad \textit{dividing by B}\\ \quad \\ ratio2\cdot A=ratio1\cdot C\implies 3(6,2)=2(8,4)\\ \quad \\\qquad {\color{blue}{ B=\left(\cfrac{\textit{sum of "x" values}}{ratio1+ratio2}\quad ,\quad \cfrac{\textit{sum of "y" values}}{ratio1+ratio2}\right)}}\\ \quad \\ \qquad thus\qquad \\ \quad \\ B=\left(\cfrac{(3\cdot 6)+(2\cdot 8)}{2+3}\quad ,\quad \cfrac{(3\cdot 2)+(2\cdot 4)}{2+3}\right)\)

OpenStudy (anonymous):

Ok thanks

OpenStudy (jdoe0001):

yw